A meter rule is hung by two light strip tied at the

15cm and 75cm mark respectively to a ceiling. If the

weight of the ruler is 2N. (i) calculate the tension in

tge two string.(ii) calculate also the force that must

be applied at the 100cm mark for the first string slack


1
Expert's answer
2021-09-03T13:56:27-0400
"\sum M_A=0:\quad 2\:\rm N*0.35\: m-R_B*0.6\: m=0"

"R_B=1.2\:\rm N"

"\sum F_y=0:\quad R_A-2\:\rm N+1.2\: N=0"

"R_A=0.8\:\rm N"

"\sum M_B=0:\quad -2\:\rm N*0.25\: m+F*0.25\: m=0"

"F=2\:\rm N"